Solve each equation for y. Then find the value of y for each value of x.
x = 9 – 3y; x = –3, 6, 12

Answer by macston(5194)

.
x=9-3y
x-9=-3y
9-x=3y
(9-x)/3=y
3-(x/3)=y
ANSWER: y=3-(x/3)
.
For x=-3:
y=3-(-3/3)=3+1=4
ANSWER:(-3,4)
.
For x=6:
y=3-(6/3)=3-2=1
ANSWER: (6,1)
.
Try the last one on your own.